OpenStudy (anonymous):

____ log2 16 √ 8 =log 2 2^4*2^3/2 = log 2 2^11/2 how do you get the 11/2?? wouldnt it b 12/2?

OpenStudy (anonymous):

is that log base 2 ?

OpenStudy (callisto):

\[\log_{2} 16\sqrt{8} = \log_{2} 2^{4\times}2^{3/2} = (4+3/2) \log_{2} 2 = 11/2\]\[**\log_{2}2 = 1 \]

OpenStudy (anonymous):

we know base is 2... i'm not writing it cuz it create confusion log rad (16 x 16 x 8 ) 1/2log 2x2x2x2x2x2x2x2x2x2x2 1/2 log 2^11 11/2log2 11/2